### v1

# Clustering in Networks with Time-varying Nodal Attributes

Abstract
This manuscript studies nodal clustering in graphs having a time series at each node. The framework includes priors for low-dimensional representations and a decoder that bridges the latent representations and time series. The structural and temporal patterns are fused into representations that facilitate clustering, addressing the limitation that the evolution of nodal attributes is often overlooked. Parameters are learned via maximum approximate likelihood, with a graph-fused LASSO regularization imposed on prior parameters. The optimization problem is solved via alternating direction method of multipliers; Langevin dynamics are employed for posterior inference. Simulation studies on block and grid graphs with autoregressive dynamics, and applications to California county temperatures and a book word co-occurrence network demonstrate the effectiveness of the proposed method.
Keywords: ADMM, LASSO, Latent Space Models, Representation Learning, Time Series.
1 Introduction
Networks are used to describe relational phenomena, where entities are denoted by nodes and relations are delineated by edges between nodes. Such structures naturally arise in various domains, including social interactions (snijders2011statistical), neurology (yang2022functional), and finance (jackson2021systemic). A fundamental task in network analysis involves clustering nodes into disjoint sets, where node within a set are more connected than nodes in disjoint sets. Such clusters often provide structural insights into the network (see shai2017case for a collection of case studies). The majority of clustering methods focus on structural patterns (abbe2018community). However, many networks possess additional “nodal attributes”. This paper shows how to take advantage of nodal time series, providing clustering methods that account for both network structure and attribute dynamics. Section 4 uses the methods to identify climate regions in California by county. Here, both spatial proximities and temperatures of the counties are used to develop a realistic clustering.
Clustering nodes of networks with nodal attributes relies on effectively combining graphical structures with the nodal information. Methods have been previously proposed for clustering attributed networks. binkiewicz2017casc perform covariate-assisted spectral clustering by augmenting the Laplacian of the graph with pairwise similarities of the nodal covariates. shen2024bayesian propose a Bayesian stochastic block model where nodal information is incorporated into prior distributions via a covariate-dependent random partition prior. hu2024nac leverage individual and neighboring covariates with a node-specific weight matrix to improve clustering. Furthermore, the graph-fused LASSO regularizations in chen2023more and Padilla2023graphon have shown promising results, producing smoothed signals between neighboring nodes while preserving boundaries between regions having distinct signals. Following hallac2015network and yu2025graph, our framework also adapts graph-fused LASSO regularization to achieve similar filtering behavior.
In many applications, nodal attributes progress in time, while the network structure remains fixed. Nodal time series can provide valuable clustering information, especially in cases with a non-informative graph. Some methods for clustering graphs nodes overlook the evolution of nodal features, eschewing an approach that merges temporal dynamics with structural patterns. A natural solution introduces a latent variable at each node. By compressing each nodal series into a low-dimensional latent representation, temporal dynamics can be summarized and integrated with structural patterns. In our approach, latent representations are managed by node-specific priors and empirical Bayes techniques. Simultaneously, the latent space is regularized by a graph-fused LASSO penalty to incorporate network information. The resulting low-dimensional representations preserve both neighbor information and attribute dynamics, promoting effective clustering.
Latent space models bridge low-dimensional representations with the observed nodal series. handcock2007model model nodes in a Euclidean space by assuming that edge probabilities are greater for node pairs that are closer in the latent space. zhu2023disentangling construct a latent space model with both attractive and repulsive nodes, thereby disentangling network polarization dynamics. In contrast, our framework models the nodal series, conditioning on its low-dimensional representations. Unlike a variational auto-encoder (kingma2013auto), which encourages the approximate posterior to align with a zero-mean Gaussian prior, we focus on learning Gaussian prior means to facilitate clustering. With graph-fused LASSO regularization imposed on the differences between prior parameters over the edges, nodes having similar patterns are encouraged to have similar prior parameters. Consequently, a $k$-means algorithm that clusters via centroids is well-suited.
This manuscript makes the following contributions:
  • A decoder-only latent space model is proposed to link the nodal series and latent variables. Anchored on the decoder, the latent variables are regarded as low-dimensional representations of the nodal series, while also capturing network structural patterns. The prior parameters are clustered via $k$-means method.
  • We show how to learn the Gaussian prior parameters in our latent representation. Specifically, our graph-fused LASSO regularization encourages sparsity in the multivariate parameters across edges. This produces similar parameters in neighboring nodes while preserving regions with disparate signals.
  • An alternating direction method of multipliers (ADMM) procedure is developed to solve our associated optimization problem; parameters are learned via posterior inference. Through simulations on block and grid graphs with autoregressive dynamics, and applications to California county temperatures and a word co-occurrence network, the effectiveness of the proposed framework is demonstrated.
The rest of the manuscript is organized as follows. Section 2 describes our proposed framework. Section 3 presents an objective function with a graph-fused LASSO regularization and develops the ADMM procedure to solve the optimization problem. Section 4 illustrates the proposed method through simulations and real-data applications. Section 5 discusses limitations and ideas for future work.
2 Decoder-only Latent Space Models
We consider an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the node set and $\mathcal{E}$ is the edge set. For each node $i\in\mathcal{V}$, let $\{Y_{t,i}\}_{t=1}^{n}$ denote a sequence of univariate random variables (the node $i$ time series). The node $i$ series is represented as $\bm{Y}_{i}\coloneqq(Y_{1,i},Y_{2,i},\dots,Y_{n,i})^{\top}\in\mathbb{R}^{n}$. For each $\bm{Y}_{i}\in\mathbb{R}^{n}$, we assume existence of a latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ such that $\bm{Y}_{i}$ is generated from $\bm{Z}_{i}$ with the following time series “decoder”:
$\bm{Y}_{i}|\bm{Z}_{i}\sim P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})\stackrel{{\scriptstyle\cal D}}{{=}}\mathcal{N}\big(\bm{h}_{\bm{\phi}}(\bm{Z}_{i}),\bm{I}_{n}\big),$
where $\bm{h}_{\bm{\phi}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ is a neural network parameterized by $\bm{\phi}$. Here, $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ is the conditional probability density function of $\bm{Y}_{i}$ given $\bm{Z}_{i}$ (which depends on the neural network parameters in $\bm{\phi}$). We impose a prior distribution on the latent $\bm{Z}_{i}$ via
$\bm{Z}_{i}\sim P_{\bm{\mu}_{i}}(\bm{Z}_{i})\stackrel{{\scriptstyle\cal D}}{{=}}\mathcal{N}(\bm{\mu}_{i},\bm{I}_{d}).$
The prior parameter, $\bm{\mu}_{i}\in\mathbb{R}^{d}$, is learned for each node $i\in\mathcal{V}$. For the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$, the latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ is a low-dimensional representation of $\bm{Y}_{i}\in\mathbb{R}^{n}$ at node $i$.
Here, the distribution of interest is $P(\bm{Y}_{i})$, which may be complicated. Instead of directly imposing a parametric form on the marginal $P(\bm{Y}_{i})$, our framework introduces an auxiliary latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ to model $P(\bm{Y}_{i})$ as a mixture:
$P(\bm{Y}_{i})=\int P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})P_{\bm{\mu}_{i}}(\bm{Z}_{i})d\bm{Z}_{i}.$
While the flexibility of our framework allows $\bm{Y}_{i}$ to be a vector of multivariate features, we regard $\bm{Y}_{i}\in\mathbb{R}^{n}$ as a univariate series of length $n$. Two remarks follow.
Remark 1.
The Gaussian assumption is placed on $P(\bm{Y}_{i}|\bm{Z}_{i})$, not on the marginal $P(\bm{Y}_{i})$. The induced distribution of $P(\bm{Y}_{i})$ can be highly non-Gaussian and depends on $\bm{h}_{\bm{\phi}}(\bm{\cdot})$.
Remark 2.
Our framework assumes that $Y_{t,i}$ is independent in time conditional on $\bm{Z}_{i}$. Here, $\bm{Z}_{i}$ serves as our low-dimensional representation that captures temporal relationships while retaining a flexible marginal for $P(\bm{Y}_{i})$. In particular, the covariance of $\bm{Y}_{i}$ is
$\mathrm{Cov}(\bm{Y}_{i})$ $=\mathbb{E}\big[\mathrm{Cov}(\bm{Y}_{i}|\bm{Z}_{i})\big]+\mathrm{Cov}\big[\mathbb{E}(\bm{Y}_{i}|\bm{Z}_{i})\big]=\bm{I}_{n}+\mathrm{Cov}\big[\bm{h}_{\bm{\phi}}(\bm{Z}_{i})\big]$
which depends on $\bm{h}_{\bm{\phi}}(\bm{\cdot})$, thereby allowing for marginal temporal dependence.
For flexibility, the mean of $P(\bm{Y}_{i}|\bm{Z}_{i})$ is modeled with neural networks, which have universal approximation properties (hornik1991approximation; yarotsky2017error). Specifically, $\bm{h}_{\bm{\phi}}(\bm{\cdot})$ is taken as a three-layer rectified linear unit (ReLU) neural network:
$\bm{h}_{\bm{\phi}}(\bm{Z}_{i})=\bm{W}_{3}\mathrm{ReLU}\big(\bm{W}_{2}\mathrm{ReLU}(\bm{W}_{1}\bm{Z}_{i}+\bm{b}_{1})+\bm{b}_{2}\big)+\bm{b}_{3},$
where $\bm{W}_{1}\in\mathbb{R}^{h_{1}\times d},\bm{W}_{2}\in\mathbb{R}^{h_{2}\times h_{1}}$, and $\bm{W}_{3}\in\mathbb{R}^{n\times h_{2}}$ are weights and $\bm{b}_{1}\in\mathbb{R}^{h_{1}\times 1},\bm{b}_{2}\in\mathbb{R}^{h_{2}\times 1}$, and $\bm{b}_{3}\in\mathbb{R}^{n\times 1}$ are bias parameters. Here, $\mathrm{ReLU}(x)=\max(0,x)$ is the element-wise ReLU activation function at $x$, and model parameters are collected into $\bm{\phi}$. While other choices of depth, activation functions, and architectures (lecun2015deep) can be used, ReLU neural networks are commonly employed and have been extensively studied.
Figure 1 overviews our framework. The shaded circles in the top layer depict the series $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$, and the dashed circle in the bottom layer depicts the latent $\{\bm{Z}_{i}\}_{i\in\mathcal{V}}$. The series are produced from the latent variables in a bottom-up manner. The decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ with neural network parameter $\bm{\phi}$ is shared across all nodes, while the means $\bm{h}_{\bm{\phi}}(\bm{Z}_{i})$ differ by nodes. Intuitively, the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ helps learn the time series representation $\bm{Z}_{i}$ in a top-down manner.
$\bm{Z}_{1}$$\bm{Y}_{1}$$\mathcal{N}(\bm{\mu}_{1},\bm{I}_{d})$$\bm{Z}_{2}$$\bm{Y}_{2}$$\mathcal{N}(\bm{\mu}_{2},\bm{I}_{d})$$\bm{Z}_{3}$$\bm{Y}_{3}$$\mathcal{N}(\bm{\mu}_{3},\bm{I}_{d})$$\bm{Z}_{4}$$\bm{Y}_{4}$$\mathcal{N}(\bm{\mu}_{4},\bm{I}_{d})$$\bm{Z}_{5}$$\bm{Y}_{5}$$\mathcal{N}(\bm{\mu}_{5},\bm{I}_{d})$$P_{\bm{\phi}}(\bm{Y}_{1}|\bm{Z}_{1})$$P_{\bm{\phi}}(\bm{Y}_{2}|\bm{Z}_{2})$$P_{\bm{\phi}}(\bm{Y}_{3}|\bm{Z}_{3})$$P_{\bm{\phi}}(\bm{Y}_{4}|\bm{Z}_{4})$$P_{\bm{\phi}}(\bm{Y}_{5}|\bm{Z}_{5})$
Figure 1: An illustration of our prior distributions and a decoder on a graph with five nodes.
3 Learning and Inference
This section develops methods to learn the graph and series structures. A penalized likelihood is formulated first; the resulting constrained optimization is then solved with an ADMM algorithm. Denote the joint log-likelihood of $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ as $L(\bm{\phi},\bm{\mu})=\sum_{i\in\mathcal{V}}\ell(\bm{\phi},\bm{\mu}_{i})$, where the node $i$ marginal log-likelihood is
$\ell(\bm{\phi},\bm{\mu}_{i})=\ln\big(P(\bm{Y}_{i};\bm{\phi},\bm{\mu}_{i})\big)=\ln\left(\int P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})P_{\bm{\mu}_{i}}(\bm{Z}_{i})d\bm{Z}_{i}\right).$
To facilitate clustering of the prior parameters, $\hat{\bm{\mu}}$, the penalized optimization
$\hat{\bm{\phi}},\hat{\bm{\mu}}=\operatorname*{arg\,min}_{\bm{\phi},\bm{\mu}}\left[-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\mu}_{i}-\bm{\mu}_{j}\|_{2}\right]$ (1)
is considered, where $\lambda>0$ is a tuning parameter for the penalty term. The $i$th row of the $\mathbb{R}^{N\times d}$ matrix $\bm{\mu}$ is denoted by $\bm{\mu}_{i}\in\mathbb{R}^{d}$, where $N\coloneqq|\mathcal{V}|$.
The graph-fused LASSO regularization manages associations between nodes, incorporating structural information into our latent representations. The regularization here is a total variation sum of the multivariate pairwise differences $\bm{\mu}_{i}-\bm{\mu}_{j}\in\mathbb{R}^{d}$ over edges $(i,j)\in\mathcal{E}$. Our regularization induces sparsity in the differences while permitting simultaneous shifts across nodes. By penalizing the total variation of the prior parameters across edges, our framework can identify meaningful boundaries between clusters having disparate signals.
To optimize (1) (which involves latent variables), we first manipulate the objective function. Introduce the slack variables $\bm{\nu}_{i,j}\in\mathbb{R}^{d}$ and recast the constrained optimization as
$\hat{\bm{\phi}},\hat{\bm{\mu}}=\operatorname*{arg\,min}_{\bm{\phi},\bm{\mu}}\left[-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}\right],\qquad\text{subject to}\qquad\bm{\mu}_{i}-\bm{\mu}_{j}=\bm{\nu}_{i,j}.$ (2)
The augmented Lagrangian is
$\mathcal{L}(\bm{\phi},\bm{\mu},\bm{\nu},\bm{\rho})=-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}+\sum_{(i,j)\in\mathcal{E}}\bm{\rho}_{i,j}^{\top}(\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j})+\frac{\gamma}{2}\sum_{(i,j)\in\mathcal{E}}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}\rVert_{2}^{2},$
where $\bm{\rho}_{i,j}\in\mathbb{R}^{d}$ is the Lagrange multiplier for each $(i,j)\in\mathcal{E}$ and $\gamma>0$ is an additional penalty parameter for the augmentation term. Let $\bm{w}_{i,j}=\gamma^{-1}\bm{\rho}_{i,j}\in\mathbb{R}^{d}$ denote a scaled dual variable. The augmented Lagrangian can be equivalently expressed as
$\mathcal{L}(\bm{\phi},\bm{\mu},\bm{\nu},\bm{w})=-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}+\frac{\gamma}{2}\sum_{(i,j)\in\mathcal{E}}\left(\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j}\rVert_{2}^{2}-\lVert\bm{w}_{i,j}\rVert_{2}^{2}\right).$
An ADMM procedure (boyd2011distributed) recursively solves our optimization:
$\bm{\phi}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\phi}}\left[-L(\bm{\phi},\bm{\mu}^{(a)})\right],$ (3)
$\bm{\mu}_{i}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\mu}_{i}}\left[-L(\bm{\phi}^{(a+1)},\bm{\mu}_{i})+\frac{\gamma}{2}\sum_{j\in\mathcal{B}(i)}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}^{(a)}+\bm{w}_{i,j}^{(a)}\rVert_{2}^{2}\right],\qquad\forall\ i\in\mathcal{V},$ (4)
$\bm{\nu}_{i,j}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\nu}_{i,j}}\left[\lambda\|\bm{\nu}_{i,j}\|_{2}+\frac{\gamma}{2}\lVert\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}+\bm{w}_{i,j}^{(a)}\rVert_{2}^{2}\right],\qquad\forall\ (i,j)\in\mathcal{E},$ (5)
$\bm{w}_{i,j}^{(a+1)}$ $=\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}^{(a+1)}+\bm{w}_{i,j}^{(a)},\qquad\forall\ (i,j)\in\mathcal{E}.$ (6)
Here, $a$ denotes the ADMM algorithm iteration and $\mathcal{B}(i)\coloneqq\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$ is the set of node $i$’s neighbors.
The solution $\bm{\mu}_{i}$ at a particular iteration of the ADMM procedure is
$\bm{\mu}_{i}=\left(1+\gamma|\mathcal{B}(i)|\right)^{-1}\left[\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})+\gamma\sum_{j\in\mathcal{B}(i)}(\bm{\mu}_{j}+\bm{\nu}_{i,j}-\bm{w}_{i,j})\right].$ (7)
Moreover, the gradient with respect to the decoder parameter is
$\nabla_{\bm{\phi}}\ L(\bm{\phi},\bm{\mu})=-\sum_{i\in\mathcal{V}}\mathbb{E}\Big(\nabla_{\bm{\phi}}\ln\big(P(\bm{Y}_{i}|\bm{Z}_{i})\big)\Big|\bm{Y}_{i}\Big).$ (8)
The parameter $\bm{\phi}$ is updated via back-propagation. The update in (5) is equivalent to solving a group LASSO problem (yuan2006model; alaiz2013group). One can iteratively update $\bm{\nu}_{i,j}$ for each edge $(i,j)\in\mathcal{E}$ via
$\bm{\nu}_{i,j}=\left(1-\frac{\lambda}{\gamma\|\bm{s}_{i,j}\|_{2}}\right)_{+}\bm{s}_{i,j},$ (9)
where $\bm{s}_{i,j}=\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}$ and $(x)_{+}=\max(0,x)$.
Calculating (7) and the gradient in (8) requires evaluating a conditional expectation under the posterior density $P(\bm{Z}_{i}|\bm{Y}_{i})\propto P(\bm{Y}_{i}|\bm{Z}_{i})\times P(\bm{Z}_{i})$. Here, Langevin dynamics are employed to sample from the posterior distribution, approximating conditional expectations (du2019implicit; nijkamp2020anatomy; purohit2024posterior). In particular, let $k$ be the time step of the Langevin dynamics and let $\delta>0$ be a small step size. The Langevin dynamics to draw samples from the posterior distribution iterate
$\bm{Z}_{i}^{k+1}$ $=\bm{Z}_{i}^{k}+\delta\big[\nabla_{\bm{Z}_{i}}\ln\big(P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i}^{k})\big)-(\bm{Z}_{i}^{k}-\bm{\mu}_{i})\big]+\sqrt{2\delta}\bm{\epsilon}$ (10)
in $k$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ is a random perturbation. The gradient of $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ with respect to $\bm{Z}_{i}$ can be calculated through back-propagation. The derivations and computation complexity of the ADMM procedure are provided in our Supplementary Material.
3.1 Nodal Clusterings
After parameter learning, $k$-means method is applied to $\{\hat{\bm{\mu}}_{i}\}$ to cluster the nodes. The LASSO regularization of $\sum_{(i,j)\in\mathcal{E}}\|\bm{\mu}_{i}-\bm{\mu}_{j}\|_{2}$ encourages connected nodes to have similar prior parameters. As a result, the estimated $\{\hat{\bm{\mu}}_{i}\}$ naturally “cluster well” in Euclidean spaces, making the $k$-means method well-suited for partitioning.
The nodal series may contain temporal similarity that are shared by neighboring nodes. By jointly modeling temporal dynamics and structural patterns through the prior parameters, our framework integrates both sources of information to form clusters. Consequently, the resulting partitions are more informative, in the sense that they reflect both temporal similarity and structural proximity, rather than solely relying on the time series $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ or the graph structure $\mathcal{G}=(\mathcal{V},\mathcal{E})$ alone.
Since most $k$-means algorithms require the number of clusters, $k$, we estimate this parameter via a silhouette score. Specifically, a range of $k$ in $\{2,\dots,K\}$ is considered; the one having the highest silhouette score $S\in[-1,1]$, defined by
$S=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\frac{d_{i}^{\text{out}}-d_{i}^{\text{in}}}{\max(d_{i}^{\text{in}},d_{i}^{\text{out}})},$
is selected. Here, $d_{i}^{\text{in}}$ is the mean distance between point $i$ and all points in the same cluster, and $d_{i}^{\text{out}}$ is the minimum mean distance between point $i$ and any other cluster. A higher $S$ suggests more cohesive clusters, with distinct separation between different clusters.
4 Experiments
4.1 Simulations
To generate realistic networks, block and grid graphs will be simulated. Block graphs mimic social relations, where individuals with similar attributes often form cohesive clusters. Grid graphs imitate spatial aspects such as urban layouts or power systems, where entities reside on a lattice and clusters reflect spatial regions. To simulate realistic cluster-specific time series, autoregressive (AR) and vector autoregressive (VAR) models are used. Both models inject temporal dependence by linearly regressing on past series values.
Our simulations generate $50$ graph realizations with a varying number of nodes $N$. The series length $n=100$ is used at every node. To evaluate performance, the following standard clustering metrics are considered: normalized mutual information (NMI), adjusted Rand index (ARI), accuracy score (ACC), homogeneity (HOM), completeness (COM), and cluster purity (PUR). These evaluation metrics require a known truth to measure from; see emmons2016analysis and su2022comprehensive for additional detail. Six competitors, $k$-means, dynamic time warping (DTW) (sakoe2003dtw), covariate assisted spectral clustering (CASC) (binkiewicz2017casc), semi-definite programming (SDP) (yan2021sdp), spectral clustering on network-adjusted covariates (NAC) (hu2024nac), and spectral clustering on ratios-of-eigenvectors (SCORE) (Jin2015SCORE), are provided for comparison.
The $k$-means and DTW methods only use the series for clustering, while the SCORE method only uses graphical structure. The CASC, SDP, and NAC are hybrid methods utilizing both series and graphical structures to cluster nodes, similar to our method. Our simulations assess whether integrating both sources of information leads to better clustering than approaches that use the graph or attributes alone. Performances between our method and existing hybrid approaches are compared. Details of the competitors are provided in the Supplementary Material.
Throughout, our method is dubbed GFL. The latent dimension is taken as $d=3$, and the two hidden layers of the neural network each contain $32$ neurons. The tuning $\lambda$’s are $\{0.1,0.25,0.5,0.75,1.0\}$ and $10\%$ of the series was held out for model selection. For each $\lambda$, $50$ iterations of the ADMM procedure are run and $\gamma=\lambda$ is taken. In each ADMM iteration, the neural network is trained using $20$ steps of an Adam optimizer with a learning rate of $10^{-4}$, while $20$ iterations of block coordinate descent are run for group LASSO. For Langevin dynamic sampling, we take $\delta=0.4$ and $500$ samples are drawn for each node using $50$ MCMC iterations. As the objective function in (2) is highly non-convex due to the neural networks and sub-routines requiring MCMC sampling, specifying standard convergence criteria is impractical. The algorithm is implemented for a preset number of iterations, as an early stopping strategies to avoid over-fitting, which is commonly used in deep learning. The model selection procedure and results for varying latent dimensions as a sensitivity analysis are presented in our Supplementary Material.
Scenario 1
Our first scenario simulates graphs with block structures to demonstrate nodal clustering. Specifically, for the node set $\mathcal{V}$ of size $N=|\mathcal{V}|$, the probability matrix $\bm{P}\in[0,1]^{N\times N}$ is defined with entries ${\bm{P}}_{i,j}=0.30\ \forall\ i,j\in\mathcal{C}_{k},\ k\in\{1,2,3\}$, and ${\bm{P}}_{i,j}=0.15$ otherwise. Here, $\mathcal{C}_{1},\mathcal{C}_{2}$, and $\mathcal{C}_{3}$ are three clusters that partition $\mathcal{V}$. Then edges are sampled via an adjacency matrix $\bm{A}\in\{0,1\}^{N\times N}$ format via $\bm{A}_{i,j}\sim\text{Bernoulli}({\bm{P}}_{i,j})$. For $N=120$ nodes, our three unbalanced cluster sizes are $30$, $40$ and $50$; for $N=210$, the cluster sizes are $60$, $70$, and $80$.
For node $i$ in cluster $\mathcal{C}_{k}$, its time series $\bm{Y}_{i}=\{Y_{t,i}\}_{t=1}^{T}$ is generated via a cluster-specific AR model of order one:
$Y_{t,i}=\mu_{k}+\psi_{k}(Y_{t-1,i}-\mu_{k})+\varepsilon_{t,i},\hskip 18.49988ptt=2,\ldots,T,\qquad i\in\mathcal{C}_{k},$
where $\mu_{k}$ is the mean of cluster $\mathcal{C}_{k}$, $\psi_{k}$ is the AR coefficient with $|\psi_{k}|<1$, and $\varepsilon_{t,i}\sim\mathcal{N}(0,1)$ is Gaussian white noise. The process is rendered stationary by taking the initial state
$Y_{1,i}\sim\mathcal{N}\left(\mu_{k},\frac{\sigma_{k}^{2}}{1-\psi_{k}^{2}}\right),\hskip 18.49988pti\in\mathcal{C}_{k}.$ (11)
We choose $\psi_{k}=0.5$ and $\sigma_{k}^{2}=1$ for all three clusters. Cluster means are $\mu_{1}=-1.0$, $\mu_{2}=0.0$, and $\mu_{3}=1.0$. Figure 2 depicts a simulated block graph and its associated series sample means. Table 1 reports sample means and standard deviations of our clustering evaluation metrics over the 50 simulations.
Figure 2: A simulated graph and its time series sample means. The left plot is a simulated graph with unbalanced cluster sizes and $120$ nodes. The right plot shows time series sample means along with the $\pm 1$ standard deviation bands at each time point for the three clusters.
With 120 nodes, $k$-means applied directly to the series performs reasonably well. This is because the clusters are well separated via their series means. The graph-based methods, such as CASC, SDP, NAC, and SCORE, do not perform as well, which is attributed to small graph sizes and unbalanced cluster sizes. As $N$ increases from $120$ to $210$, the performance of these methods improve, which is expected since more data is available. The DTW method also improves, and the graph-based methods benefit from a larger adjacency matrix in their spectral clusterings. Importantly, our proposed GFL method outperforms all competitor methods, highlighting its utility.
Table 1: Sample means (one standard deviation) of our evaluation metrics for Scenario 1. The best metric is in bold.
NN Method NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ HOM $\uparrow$ COM $\uparrow$ PUR $\uparrow$
$120$ GFLd=3 $\bm{98.52}\%(0.03)$ $\bm{98.96}\%(0.02)$ $\bm{99.63}\%(0.01)$ $\bm{98.49}\%(0.03)$ $\bm{98.56}\%(0.03)$ $\bm{99.63}\%(0.01)$
$k$-means $94.88\%(0.09)$ $95.25\%(0.14)$ $97.13\%(0.14)$ $96.16\%(0.04)$ $94.57\%(0.11)$ $97.13\%(0.14)$
DTW $59.76\%(0.15)$ $54.02\%(0.21)$ $69.52\%(0.19)$ $56.97\%(0.18)$ $64.98\%(0.14)$ $85.20\%(0.14)$
CASC $27.56\%(0.10)$ $26.45\%(0.12)$ $47.67\%(0.14)$ $28.99\%(0.14)$ $27.12\%(0.10)$ $61.27\%(0.11)$
SDP $37.85\%(0.08)$ $31.39\%(0.10)$ $51.27\%(0.15)$ $39.36\%(0.12)$ $37.31\%(0.08)$ $62.97\%(0.10)$
NAC $37.35\%(0.08)$ $35.86\%(0.11)$ $64.93\%(0.14)$ $38.86\%(0.12)$ $36.83\%(0.08)$ $67.67\%(0.12)$
SCORE $36.19\%(0.11)$ $37.78\%(0.15)$ $66.58\%(0.17)$ $37.53\%(0.14)$ $35.82\%(0.11)$ $70.40\%(0.13)$
$210$ GFLd=3 $\bm{98.91}\%(0.02)$ $\bm{99.29}\%(0.01)$ $\bm{99.69}\%(0.01)$ $\bm{99.11}\%(0.01)$ $\bm{98.73}\%(0.02)$ $\bm{99.69}\%(0.01)$
$k$-means $95.39\%(0.09)$ $95.86\%(0.14)$ $97.30\%(0.14)$ $96.71\%(0.03)$ $95.12\%(0.11)$ $97.30\%(0.14)$
DTW $79.11\%(0.09)$ $80.64\%(0.14)$ $91.46\%(0.14)$ $80.22\%(0.08)$ $79.08\%(0.11)$ $91.83\%(0.13)$
CASC $81.76\%(0.09)$ $85.76\%(0.13)$ $93.72\%(0.13)$ $83.11\%(0.06)$ $81.46\%(0.11)$ $93.72\%(0.13)$
SDP $39.17\%(0.05)$ $35.18\%(0.08)$ $60.02\%(0.13)$ $40.51\%(0.10)$ $38.87\%(0.06)$ $64.43\%(0.10)$
NAC $42.85\%(0.06)$ $42.04\%(0.10)$ $70.17\%(0.14)$ $44.19\%(0.10)$ $42.57\%(0.07)$ $72.11\%(0.12)$
SCORE $75.49\%(0.09)$ $80.38\%(0.13)$ $91.76\%(0.13)$ $76.75\%(0.07)$ $75.28\%(0.10)$ $91.76\%(0.13)$
Scenario 2
This scenario moves to graphs on grids having four unbalanced regional clusters: top left $\mathcal{C}_{1}$, top right $\mathcal{C}_{2}$, bottom left $\mathcal{C}_{3}$, and bottom right $\mathcal{C}_{4}$. The nodal series use the same AR model in Scenario 1 and the same covariance parameters. To incorporate regional heterogeneity, cluster means were changed to $\mu_{1}=-0.8$, $\mu_{2}=0.0$, $\mu_{3}=0.8$, and $\mu_{4}=1.6$.
We consider two graph sizes having $N=144$ and $N=196$ nodes, corresponding to $12\times 12$ and $14\times 14$ grids. Both densities are sparser than Scenario 1. The unbalanced cluster sizes are $25$, $30$, $42$, and $47$ for $N=144$ and $35$, $40$, $42$, and $79$ for $N=196$. Figure 3 illustrates a simulated grid and its associated time series means. No clusters of nodes are visually distinguishable from the graph topology (without relying on nodal colors). Table 2 reports the sample means and standard deviations of the evaluation metrics across 50 simulations.
Figure 3: A simulated grid and its associated time series means. The left plot is a simulated graph with 144 nodes having unbalanced cluster sizes. The right plot shows the sample series means of the four clusters; one standard deviation bands are again displayed.
Different from Scenario 1, the grid structure does not reveal cluster boundaries, making clustering based on the graph topology challenging. In this setting, cluster information primarily resides in the nodal time series. The hybrid methods, CASC, SDP, and NAC, which rely on both adjacency matrix and nodal series perform poorly, suggesting that an uninformative spatial structure hinders spectral clustering even when the time series exhibit cluster differences. Moreover, increasing the graph size from $N=144$ to $N=196$ does not improve their performances, and can even degrade matters. Enlarging the grid jeopardizes graph-based approaches that emphasize structural cues. Interestingly, the SCORE method, which does not utilize any series for clustering, performs well when the correct number of clusters is supplied. The slight error here stems from the method’s tendency to produce four equal-sized clusters, making some boundary nodes incorrectly assigned. In contrast, our GFL method continues to perform nicely, adapting to the absence of cluster signals in graph and relying on informative nodal series.
Table 2: Sample means (one standard deviation) of our evaluation metrics for Scenario 2. The best metric is in bold.
NN Method NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ HOM $\uparrow$ COM $\uparrow$ PUR $\uparrow$
$144$ GFLd=3 $\bm{99.24}\%(0.01)$ $\bm{99.36}\%(0.01)$ $\bm{99.76}\%(0.01)$ $\bm{99.22}\%(0.02)$ $\bm{99.26}\%(0.01)$ $\bm{99.76}\%(0.01)$
$k$-means $83.87\%(0.10)$ $82.86\%(0.16)$ $90.64\%(0.17)$ $85.05\%(0.08)$ $83.53\%(0.11)$ $92.13\%(0.14)$
DTW $58.34\%(0.06)$ $46.60\%(0.11)$ $65.82\%(0.16)$ $58.37\%(0.09)$ $59.29\%(0.08)$ $75.89\%(0.12)$
CASC $32.68\%(0.03)$ $22.32\%(0.04)$ $31.25\%(0.06)$ $34.11\%(0.10)$ $32.10\%(0.02)$ $46.76\%(0.06)$
SDP $28.92\%(0.05)$ $18.43\%(0.04)$ $36.50\%(0.10)$ $30.30\%(0.11)$ $28.37\%(0.05)$ $50.26\%(0.07)$
NAC $32.79\%(0.09)$ $23.45\%(0.09)$ $47.32\%(0.12)$ $34.08\%(0.13)$ $32.32\%(0.09)$ $56.15\%(0.10)$
SCORE $74.95\%(0.05)$ $69.58\%(0.10)$ $86.49\%(0.12)$ $76.93\%(0.03)$ $73.81\%(0.07)$ $86.49\%(0.12)$
$196$ GFLd=3 $\bm{99.71}\%(0.01)$ $\bm{99.75}\%(0.01)$ $\bm{99.92}\%(0.01)$ $\bm{99.71}\%(0.01)$ $\bm{99.71}\%(0.01)$ $\bm{99.92}\%(0.01)$
$k$-means $85.48\%(0.08)$ $86.10\%(0.13)$ $93.52\%(0.13)$ $86.98\%(0.04)$ $84.87\%(0.09)$ $93.52\%(0.13)$
DTW $57.66\%(0.07)$ $47.02\%(0.12)$ $65.21\%(0.17)$ $58.24\%(0.09)$ $58.08\%(0.08)$ $74.69\%(0.12)$
CASC $28.53\%(0.02)$ $20.30\%(0.03)$ $28.91\%(0.04)$ $30.29\%(0.10)$ $27.70\%(0.01)$ $44.85\%(0.06)$
SDP $27.22\%(0.09)$ $17.94\%(0.07)$ $28.67\%(0.07)$ $28.92\%(0.13)$ $26.43\%(0.08)$ $48.37\%(0.08)$
NAC $30.42\%(0.08)$ $21.43\%(0.07)$ $40.19\%(0.12)$ $32.08\%(0.12)$ $29.67\%(0.07)$ $53.61\%(0.09)$
SCORE $66.17\%(0.04)$ $55.89\%(0.08)$ $80.54\%(0.11)$ $68.79\%(0.04)$ $64.50\%(0.06)$ $80.54\%(0.11)$
Scenario 3
Scenario 3 extends Scenario 1 by including correlation within series in the same cluster. Here, a VAR model of order one is applied to incorporate cluster-specific correlations. The $N$-dimensional time series $\{{{\bm{Y}}_{t}}\}_{t=1}^{T}$ are generated simultaneously over all nodes via
${\bm{Y}}_{t}={\bm{\beta}}+{\Phi}({\bm{Y}}_{t-1}-{\bm{\beta}})+{\bm{\xi}}_{t},\qquad t=2,\ldots,T$
where $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{N})^{\top}$ is the nodal mean vector: $\beta_{i}=\mu_{k}$ for $i\in\mathcal{C}_{k}$. Here, the error $\bm{\xi}_{t}$ is sampled as $\mathcal{N}({\bm{0}}_{N},\bm{\Sigma})$ for each $t$, where $\bm{\Sigma}=\{\Sigma_{i,j}\}_{i,j=1,\ldots,N}$ has the intra-cluster correlations $\Sigma_{i,j}=1\ \forall\ i=j$, $\Sigma_{i,j}=\rho\ \forall\ i\neq j,\text{ and }i,j\in\mathcal{C}_{k},\ k\in\{1,2,3\}$, and $\Sigma_{i,j}=0$ otherwise. We take $\rho=0.3$, which induces moderate correlation.
The VAR coefficient is taken as $\Phi=0.5\bm{I}_{N}$ and the means for the three clusters are repeated from Scenario 1: $\mu_{1}=-1.0$, $\mu_{2}=0.0$, and $\mu_{3}=1.0$. Series from different clusters are statistically independent. Our simulations start with ${\bm{Y}}_{1}={\bm{0}}$ and burn-in is used for $100$ iterations to ensure stationary nodal series. The same three cluster block graph in Scenario 1 is used. Figure 4 shows a simulated grid graph and the associated series sample means. Comparing to Scenario 1, the series means are not as “well separated”, making clustering more difficult.
Table 3 reports the means and standard deviations of our evaluation metrics over 50 simulations. With $120$ nodes, $k$-means and CASC methods again provide good clusterings. CASC clusterings, which leverage nodal covariates via pair-wise similarity, benefits from the positive intra-cluster correlations. When the number of nodes increases to $210$, competitors improve significantly (more than in Scenario 1), exploiting the temporal correlation between nodes to improve performance. Our GFL method slightly degrades from Scenario 1, but still maintains accuracy and is often the best method. Even in the presence of modest intra-nodal correlation, our GFL clustering method continues to perform well.
Figure 4: A simulated graph and nodal time series means. The left plot shows a simulated graph with unbalanced cluster sizes and 120 nodes. The right plot shows the sample series means of the three clusters; one standard deviation bands are again displayed.
Table 3: Sample means (one standard deviation) of our evaluation metrics for Scenario 3. The best metric is in bold.
NN Method NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ HOM $\uparrow$ COM $\uparrow$ PUR $\uparrow$
[…]
$k$-means $\bm{98.73}\%(0.08)$ $98.00\%(0.14)$ $98.05\%(0.14)$ $\bm{100.0}\%(0.00)$ $99.45\%(0.11)$ $98.05\%(0.14)$
DTW $70.09\%(0.20)$ $65.29\%(0.26)$ $77.87\%(0.21)$ $68.23\%(0.22)$ $74.12\%(0.19)$ $87.87\%(0.15)$
CASC $\bm{98.73}\%(0.09)$ $98.00\%(0.14)$ $98.05\%(0.14)$ $\bm{100.0}\%(0.00)$ $98.45\%(0.11)$ $98.05\%(0.14)$
SDP $90.44\%(0.17)$ $85.92\%(0.26)$ $87.37\%(0.24)$ $91.08\%(0.17)$ $90.84\%(0.17)$ $93.67\%(0.15)$
NAC $81.64\%(0.09)$ $84.97\%(0.13)$ $93.43\%(0.13)$ $83.16\%(0.06)$ $81.10\%(0.10)$ $93.43\%(0.13)$
SCORE $39.08\%(0.12)$ $39.98\%(0.15)$ $69.05\%(0.16)$ $40.47\%(0.15)$ $38.67\%(0.13)$ $71.83\%(0.13)$
[…]
$k$-means $\bm{98.68}\%(0.09)$ $\bm{98.00}\%(0.14)$ $98.03\%(0.14)$ $\bm{100.0}\%(0.0)$ $98.41\%(0.11)$ $98.03\%(0.14)$
DTW $89.86\%(0.10)$ $90.96\%(0.14)$ $95.59\%(0.14)$ $91.13\%(0.06)$ $89.64\%(0.11)$ $95.59\%(0.14)$
[…]
SDP $97.42\%(0.12)$ $96.60\%(0.16)$ $96.94\%(0.15)$ $98.69\%(0.08)$ $97.21\%(0.13)$ $97.50\%(0.14)$
NAC $93.04\%(0.09)$ $94.32\%(0.14)$ $96.77\%(0.14)$ $94.40\%(0.03)$ $92.73\%(0.11)$ $96.77\%(0.14)$
SCORE $74.11\%(0.09)$ $79.05\%(0.13)$ $91.28\%(0.13)$ $75.44\%(0.07)$ $73.83\%(0.10)$ $91.28\%(0.13)$
4.2 California County Temperatures
We now cluster counties in the state of California via observed temperatures. There are $N=58$ counties in the state, all of which report data. Monthly average temperatures over the 14-year period January 2011 - December 2024 are studied. The data come from the National Centers for Environmental Information (download details are listed below). Recent clusterings (by station, not county) of California climate appear in abatzoglou2009classification. While we examine the 14 year record to reduce impact of trends on results, California climate trend studies appear in ladochy2007recent and walton2018assessment.
A network with 58 nodes representing all counties was generated; two counties are connected by an edge if they share some common border. The goal here is to utilize both spatial proximity of the counties and their temperature series to identify useful climate regions within the state. Some pre-processing was implemented to obtain stationary series. Specifically, with $\{Y_{t}\}_{t=1}^{n}$ denoting a county monthly series, we convert to a standardized measurement by examining
$S_{kT+\nu}\coloneqq\frac{Y_{kT+\nu}-\hat{\mu}_{\nu}}{\hat{\sigma}_{\nu}},$
where $k\in\{0,1,\ldots,n_{\rm yr}-1\}$ and $\nu\in\{1,2,\ldots,T\}$ is the calendar month of the year. Here, $n_{\rm yr}=14$ and $T=12$ is the number of calendar months. Estimates of the month $\nu$ mean $\mu_{\nu}$ and its standard deviation $\sigma_{\nu}$ are
$\hat{\mu}_{\nu}=\frac{1}{n_{\rm yr}}\sum_{k=0}^{n_{\rm yr}-1}Y_{kT+\nu},\qquad\qquad\hat{\sigma}_{\nu}^{2}=\frac{\sum_{k=0}^{n_{\rm yr}-1}(Y_{kT+\nu}-\hat{\mu}_{\nu})^{2}}{n_{\rm yr}-1}.$
This standardization removes the monthly mean and variance seasonality of temperatures, but preserves county-level variability. In this analysis, the latent dimension $d$ is taken as three. Sensitivity analyses from the simulation study, provided in the Supplementary Material, suggest that this choice is adequate. Figure 5(a) displays the results from our methods. Ten clusters were selected by silhouette scores.
(a) GFL
(b) $k$-means ($k=10$)
Figure 5: California county clustering from GFL and $k$-means.
The clustering is sensible, placing desert Southeast, Mid-Coastal, Lower Coastal, Bay Area, Sierra Nevada Mountain, San Joaquin Valley, Coastal Northwest, and Mountainous Northeast counties in distinct clusters. California has many micro-climates and clustering its counties is challenging.
For comparison, Figure 5(b) shows a $k$-means clustering, the best competitor in simulation study, with ten clusters. Ten clusters were chosen to be comparable to our above clustering (if the number of clusters is selected by silhouette scores, fewer clusters and unrealistic results arise). This clustering is not as pleasing. Specifically, Inyo County in the Owens Valley has been placed in the desert south cluster (as have the coastal counties northwest of Los Angeles). The Northeast Mountain counties now mix with the Coastal Northwest counties. With understanding of the local geography, these are implausible.
4.3 Word Co-occurrence Network
We also analyze word usage from the novel David Copperfield by Charles Dickens in 1850. newman2006finding constructed a word co-occurrence network containing $112$ common adjectives and nouns from the book. The network nodes correspond to different words, and an undirected edge between two words indicates whether they occur adjacent to one another at any sentence within the novel.
Building on this network, we extend the data by constructing temporal trajectories for each word as a time series. Specifically, as the novel has $64$ chapters, the novel is divided into $n=64$ equally spaced segments. For each segment, we record the frequency of each word, yielding a sequence $\bm{C}_{i}=(C_{1,i},\ldots,C_{n,i})^{\top}$ that tracks usage of the $i$th word. To account for large differences in frequency usages of different words, we standardize each sequence via
$Y_{t,i}=\frac{C_{t,i}-\bar{C}_{i}}{\hat{\sigma}_{i}},\hskip 18.49988ptt=1,\dots,n,\qquad i\in\mathcal{V}_{\text{word}}.$
where $\bar{C}_{i}$ and $\hat{\sigma}_{i}$ denote the respective sample mean and standard deviation of the $i$th’s word series. This standardization ensures that the analysis focuses on the relative temporal dynamics of word usage rather than count magnitudes.
Figure 6(a) displays our clustering. Node shapes indicate the true linguistic labels (noun or adjective), while colors depict estimated clusters. The results demonstrate that most nouns (⚫) fall in a cluster (Cluster 1) and most adjectives (◼) into another (Cluster 2). The other two clusters are small, each containing no more than three words (and connected only to a single node in the remaining network). In particular, Cluster 3 contains the two words “aunt” and “family”, reflecting the prominence of family as a central theme in the novel and the pivotal role of Aunt Betsey. Unlike David’s mother, whose influence fades with her early passing, Aunt Betsey becomes the loving guardian, shaping David’s values and providing unwavering support throughout his growth.
(a) GFL
(b) $k$-means ($k=2$)
Figure 6: Clusters from GFL and $k$-means methods. Node shapes indicate whether a word is an adjective or noun, while colors depict cluster members.
Figure 6(b) displays the $k$-means clustering. Based on the linguistic nature (noun or adjective) of the words, two clusters were chosen. The $k$-means method fails to discover coherent clusters, likely because it disregards the co-occurrence structure delineated by the graph. As a result, nouns and adjectives intermix in the clusters. This contrast underscores the advantage of our method.
5 Discussion
This manuscript introduced a decoder-only latent space framework for clustering nodes in networks that have an observed time series at each node. By combining node-specific Gaussian priors, graph-fused LASSO regularization, and a neural network decoder, series and network dynamics are integrated into low-dimensional representations that facilitate clustering. Simulations and applications to California county temperatures and word co-occurrence network produced good clustering performance.
Avenues for future development are apparent. Foremost, the type of nodal data could change. Multiple series, functional data, or even a simple multivariate attribute are feasible. Our likelihood would need modification to accommodate such scenarios. The assumption of independent attributes across differing nodes could also be relaxed. On the network front, nodal relations often have a degree of strength, perhaps a continuous value in lieu of zero-one connections (krivitsky2012exponential; wilson2017stochastic). Adapting our LASSO regularization to handle such structure would further improve clustering. Time-varying network edges can also arise (malinovskaya2023statistical). Methods to accommodate both nodal and dyadic attributes over time would extend applicability of our methods.
6 Data and Code Availability
The California county temperature data is available at https://www.ncei.noaa.gov/access/monitoring/climate-at-a-glance/statewide/time-series. The adjectives and nouns data is available at https://networks.skewed.de/net/adjnoun. The code in this paper is available at https://github.com/allenkei/GraphFL.
7 Disclosure Statement
The authors report that there are no competing interests to declare.
SUPPLEMENTARY MATERIAL
Appendix A Parameter Learning
A.1 Updating $\bm{\mu}$ and $\bm{\phi}$
In this section, we derive the updates for prior parameter $\bm{\mu}$ and decoder parameter $\bm{\phi}$. Denote the collection of parameters $\{\bm{\phi},\bm{\mu}\}$ as $\bm{\theta}$. We first calculate the gradient of the log-likelihood $L(\bm{\theta})$ with respect to $\bm{\theta}$:
$\nabla_{\bm{\theta}}\ L(\bm{\theta})$ $=\nabla_{\bm{\theta}}\sum_{i\in\mathcal{V}}\ln\big(P(\bm{Y}_{i})\big)$
$=\sum_{i\in\mathcal{V}}\frac{1}{P(\bm{Y}_{i})}\nabla_{\bm{\theta}}P(\bm{Y}_{i})$
$=\sum_{i\in\mathcal{V}}\frac{1}{P(\bm{Y}_{i})}\nabla_{\bm{\theta}}\int P(\bm{Y}_{i},\bm{Z}_{i})d\bm{Z}_{i}$
$=\sum_{i\in\mathcal{V}}\frac{1}{P(\bm{Y}_{i})}\int\nabla_{\bm{\theta}}P(\bm{Y}_{i},\bm{Z}_{i})d\bm{Z}_{i}$
$=\sum_{i\in\mathcal{V}}\frac{1}{P(\bm{Y}_{i})}\int\left[P(\bm{Y}_{i},\bm{Z}_{i})\ \frac{1}{P(\bm{Y}_{i},\bm{Z}_{i})}\right]\Big[\nabla_{\bm{\theta}}P(\bm{Y}_{i},\bm{Z}_{i})\Big]d\bm{Z}_{i}$
$=\sum_{i\in\mathcal{V}}\frac{1}{P(\bm{Y}_{i})}\int P(\bm{Y}_{i},\bm{Z}_{i})\Big[\nabla_{\bm{\theta}}\ln\big(P(\bm{Y}_{i},\bm{Z}_{i})\big)\Big]d\bm{Z}_{i}$
$=\sum_{i\in\mathcal{V}}\int\frac{P(\bm{Y}_{i},\bm{Z}_{i})}{P(\bm{Y}_{i})}\Big[\nabla_{\bm{\theta}}\ln\big(P(\bm{Y}_{i},\bm{Z}_{i})\big)\Big]d\bm{Z}_{i}$
$=\sum_{i\in\mathcal{V}}\int P(\bm{Z}_{i}|\bm{Y}_{i})\Big[\nabla_{\bm{\theta}}\ln\big(P(\bm{Y}_{i},\bm{Z}_{i})\big)\Big]d\bm{Z}_{i}$
$=\sum_{i\in\mathcal{V}}\mathbb{E}\Big(\nabla_{\bm{\theta}}\ln\big(P(\bm{Y}_{i}|\bm{Z}_{i})P(\bm{Z}_{i})\big)\Big|\bm{Y}_{i}\Big)$
$=\sum_{i\in\mathcal{V}}\mathbb{E}\Big(\nabla_{\bm{\theta}}\ln\big(P(\bm{Y}_{i}|\bm{Z}_{i})\big)\Big|\bm{Y}_{i}\Big)+\sum_{i\in\mathcal{V}}\mathbb{E}\Big(\nabla_{\bm{\theta}}\ln\big(P(\bm{Z}_{i})\big)\Big|\bm{Y}_{i}\Big).$
Note that the expectation in the gradient is now with respect to the posterior density function $P(\bm{Z}_{i}|\bm{Y}_{i})\propto P(\bm{Y}_{i}|\bm{Z}_{i})\times P(\bm{Z}_{i})$. Furthermore, denote the objective function in Equation (4) as
$\mathcal{L}(\bm{\mu}_{i})=-L(\bm{\phi},\bm{\mu}_{i})+\frac{\gamma}{2}\sum_{j\in\mathcal{B}(i)}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j}\rVert_{2}^{2}.$
The gradient of $\mathcal{L}(\bm{\mu}_{i})$ with respect to the prior parameter $\bm{\mu}_{i}\in\mathbb{R}^{d}$ at a specific node $i$ is
$\nabla_{\bm{\mu}_{i}}\ \mathcal{L}(\bm{\mu}_{i})$ $=-\mathbb{E}\Big(\nabla_{\bm{\mu}_{i}}\ln\big(P(\bm{Z}_{i})\big)\Big|\bm{Y}_{i}\Big)+\gamma\sum_{j\in\mathcal{B}(i)}(\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j})$
$=-\mathbb{E}(\bm{Z}_{i}-\bm{\mu}_{i}|\bm{Y}_{i})+\gamma|\mathcal{B}(i)|\bm{\mu}_{i}+\gamma\sum_{j\in\mathcal{B}(i)}(-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j})$
$=-\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})+\big(1+\gamma|\mathcal{B}(i)|\big)\bm{\mu}_{i}+\gamma\sum_{j\in\mathcal{B}(i)}(-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j}).$
Setting the gradient $\nabla_{\bm{\mu}_{i}}\ \mathcal{L}(\bm{\mu}_{i})$ to zeros and solve for $\bm{\mu}_{i}$, we have
$\bm{0}$ $=-\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})+\big(1+\gamma|\mathcal{B}(i)|\big)\bm{\mu}_{i}+\gamma\sum_{j\in\mathcal{B}(i)}(-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j})$
$\big(1+\gamma|\mathcal{B}(i)|\big)\bm{\mu}_{i}$ $=\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})-\gamma\sum_{j\in\mathcal{B}(i)}(-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j})$
$\bm{\mu}_{i}$ $=\big(1+\gamma|\mathcal{B}(i)|\big)^{-1}\Big[\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})+\gamma\sum_{j\in\mathcal{B}(i)}(\bm{\mu}_{j}+\bm{\nu}_{i,j}-\bm{w}_{i,j})\Big].$
Similarly, the gradient of $\mathcal{L}(\bm{\phi},\bm{\mu})=-L(\bm{\phi},\bm{\mu})$ with respect to the decoder parameter $\bm{\phi}$ is
$\nabla_{\bm{\phi}}\ \mathcal{L}(\bm{\phi},\bm{\mu})=-\sum_{i\in\mathcal{V}}\mathbb{E}\Big(\nabla_{\bm{\phi}}\ln\big(P(\bm{Y}_{i}|\bm{Z}_{i})\big)\Big|\bm{Y}_{i}\Big).$
A.2 Updating $\bm{\nu}$
In this section, we present the derivation to update $\bm{\nu}_{i,j}$ for $(i,j)\in\mathcal{E}$, which is equivalent to solving a Group LASSO problem. Denote the objective function in (5) as $\mathcal{L}(\bm{\nu}_{i,j})$. When $\bm{\nu}_{i,j}\neq\bm{0}$, the gradient of $\mathcal{L}(\bm{\nu}_{i,j})$ with respect to $\bm{\nu}_{i,j}$ is
$\nabla_{\bm{\nu}_{i,j}}\mathcal{L}(\bm{\nu}_{i,j})$ $=\lambda\frac{\bm{\nu}_{i,j}}{\|\bm{\nu}_{i,j}\|_{2}}-\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j}).$
Setting the gradient to zeros, we have
$\bm{0}$ $=\frac{\lambda}{\|\bm{\nu}_{i,j}\|_{2}}\bm{\nu}_{i,j}+\gamma\cdot\bm{\nu}_{i,j}-\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})$
$\bm{\nu}_{i,j}$ $=\left(\frac{\lambda}{\|\bm{\nu}_{i,j}\|_{2}}+\gamma\right)^{-1}\Big[\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\Big]$ (12)
which involves $\bm{\nu}_{i,j}$ on both sides. Calculating the Euclidean norm of (12) on both sides and rearrange the terms, we have
$\|\bm{\nu}_{i,j}\|_{2}$ $=\left(\frac{\lambda+\gamma\|\bm{\nu}_{i,j}\|_{2}}{\|\bm{\nu}_{i,j}\|_{2}}\right)^{-1}\left\|\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\right\|_{2}$
$\lambda+\gamma\|\bm{\nu}_{i,j}\|_{2}$ $=\left\|\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\right\|_{2}$
$\|\bm{\nu}_{i,j}\|_{2}$ $=\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}-\frac{\lambda}{\gamma}.$
Plugging the result of $\|\bm{\nu}_{i,j}\|_{2}$ back into (12), the solution of $\bm{\nu}_{i,j}$ is
$\bm{\nu}_{i,j}$ $=\left(\frac{\lambda+\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}-\lambda}{\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}-\frac{\lambda}{\gamma}}\right)^{-1}\big[\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\big]$
$=\left(\frac{\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}-\frac{\lambda}{\gamma}}{\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}}\right)^{-1}\big[\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\big]$
$=\left(1-\frac{\lambda}{\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}}\right)(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}).$
Furthermore, when $\bm{\nu}_{i,j}=\bm{0}$, a subgradient vector $\bm{b}$ of $\|\bm{\nu}_{i,j}\|_{2}$ needs to satisfy that $\|\bm{b}\|_{2}\leq 1$. Since
$\bm{0}$ $\in\lambda\bm{b}-\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}),$
$\lambda\bm{b}$ $\in\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}),$
we obtain the condition that $\bm{\nu}_{i,j}$ becomes zeros when $\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}\leq\lambda$. Therefore, we can iteratively update $\bm{\nu}_{i,j}$ for $(i,j)\in\mathcal{E}$ with the following equation:
$\bm{\nu}_{i,j}=\left(1-\frac{\lambda}{\gamma\|\bm{s}_{i,j}\|_{2}}\right)_{+}\bm{s}_{i,j}$
where $\bm{s}_{i,j}=\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}$ and $(\cdot)_{+}=\max(0,\cdot)$.
A.3 Computation Complexity
The algorithm to solve (2) via ADMM is presented in Algorithm 1. The computation complexity is at least of order $O\big(A(|\mathcal{V}|skd+B|\mathcal{V}|sn+|\mathcal{V}|sd+D|\mathcal{E}|d)\big)$; additional gradient computation time is required for neural networks in sub-routines.
Elaborating, in each of $A$ iterations of ADMM, Langevin dynamics generate $s$ samples for each $i$, each requiring $k$ MCMC steps (and has dimension of $d$). Next, the decoder parameters $\bm{\phi}$ are updated via $B$ iterations of the Adam optimizer across all nodes using the $s$ generated samples. The output of neural networks has length $n$ for each node. Then, the node-specific prior parameters $\bm{\mu}_{i}$ are updated in closed form at each iteration using the MCMC samples, which incurs a computational cost that is linear in $|\mathcal{V}|sd$. Finally, the slack variables $\bm{\nu}_{i,j}$ are updated by block coordinate descent, repeated for $D$ iterations, resulting in a complexity of $O(D|\mathcal{E}|d)$. The scaled dual variables $\bm{w}_{i,j}$ are updated in closed form and effectively do not increase computational complexity.
In summary, the ADMM procedure decomposes a complex optimization problem into smaller components, targeting the smaller components individually. While our methods are computationally demanding, substantial clustering performance is gained, as demonstrated in the simulation study.
Algorithm 1 Latent space graph-fused LASSO
1:Input: learning iterations $A,B,D$, tuning parameter $\lambda$, penalty parameter $\gamma$, learning rate $\eta$, observed data $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$, initialization $\{\bm{\phi}^{(1)},\bm{\mu}^{(1)},\bm{\nu}^{(1)},\bm{w}^{(1)}\}$
2:for $a=1,\cdots,A$ do
3:   for $i\in\mathcal{V}$ do
4:    draw samples $\bm{Z}_{1,i},\dots,\bm{Z}_{s,i}$ from $P(\bm{Z}_{i}|\bm{Y}_{i})$ according to (10)
5:   end for
6:   for $b=1,\dots,B$ do
7:    $\bm{\phi}_{(b+1)}=\mathrm{Adam}\big(\bm{\phi}_{(b)},\nabla_{\bm{\phi}}\ \mathcal{L}(\bm{\phi},\bm{\mu}),\eta\big)$
8:   end for
9:   $\bm{\mu}_{i}^{(a+1)}=\big(1+\gamma|\mathcal{B}(i)|\big)^{-1}\Big[s^{-1}\sum_{u=1}^{s}\bm{Z}_{u,i}+\gamma\sum_{j\in\mathcal{B}(i)}(\bm{\mu}_{j}^{(a)}+\bm{\nu}_{i,j}^{(a)}-\bm{w}_{i,j}^{(a)})\Big],\ \forall\ i\in\mathcal{V}$
10:   Set $\widetilde{\bm{\nu}}^{(1)}=\bm{\nu}^{(a)}$
11:   for $d=1,\dots,D$ do
12:    Let $\widetilde{\bm{\nu}}_{i,j}^{(d+1)}$ be updated according to (9) for $(i,j)\in\mathcal{E}$
13:   end for
14:   Set $\bm{\nu}^{(a+1)}=\widetilde{\bm{\nu}}^{(d+1)}$
15:   $\bm{w}_{i,j}^{(a+1)}=\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}^{(a+1)}+\bm{w}_{i,j}^{(a)},\ \forall\ (i,j)\in\mathcal{E}$
16:end for
[…]
18:Output: learned prior parameters $\hat{\bm{\mu}}$
A.4 Model Selection
The optimization problem in (2) involves the tuning parameter $\lambda$. Given a list of candidate $\lambda$ values, model selection is performed via cross validation. Specifically, the observed data is split into training and testing sets via nodes. Let $\mathcal{V}_{\text{test}}$ be the node set of the testing data. During training, we replace the node $i$ data $\bm{Y}_{i}$ with the element-wise average $\bar{\bm{Y}}_{i}$ from all neighbors. The node $i$ data is held out for testing when $i\in\mathcal{V}_{\text{test}}$.
After parameters are learned from the training data, $\hat{\bm{\mu}}$ is used for parameters at removed nodes and the decoder calculates the log-likelihood of the non-training data. The $\lambda$ with the highest log-likelihood is selected. In particular, the log-likelihoods for the testing data are calculated by
$\sum_{i\in\mathcal{V}_{\text{test}}}\ln\big(P(\bm{Y}_{i})\big)=\sum_{i\in\mathcal{V}_{\text{test}}}\ln\Big[\int P(\bm{Y}_{i}|\bm{Z}_{i})P(\bm{Z}_{i})d\bm{Z}_{i}\Big]\approx\sum_{i\in\mathcal{V}_{\text{test}}}\ln\left[\frac{1}{s}\sum_{u=1}^{s}\big[P(\bm{Y}_{i}|\bm{Z}_{u,i})\big]\right]$
where $\{\bm{Z}_{u,i}\}_{u=1}^{s}$ are samples drawn from the marginal $P_{\hat{\bm{\mu}}_{i}}(\bm{Z}_{i})$ with the learned parameter $\hat{\bm{\mu}}_{i}$ for a node $i$. Eventually, with the selected $\lambda$, we learn the model parameters again via the full data, resulting in the final estimated parameters for clustering.
Appendix B Additional Details of Experiments
For particulars, $k$-means constructs clusters by minimizing within-cluster squared distances. DTW method measures similarities between time series by aligning them via a non-linear time warping. CASC method leverages the nodal covariate $\bm{X}$ via $\bm{L}_{\tau}\bm{L}_{\tau}+\alpha\bm{X}\bm{X}^{\top}$ to improve spectral clustering performance; here, $\bm{L}_{\tau}$ is the regularized graph Laplacian. SDP method converts clustering issues into a convex semi-definite procedure, enabling efficient approximation of the optimal partition. Specifically, SDP maximizes the inner product between $\bm{A}+\lambda\bm{K}$ and a cluster label matrix; here, $\bm{A}$ is an adjacency matrix and $\bm{K}\in\mathbb{R}_{+}^{N\times N}$ is a kernel matrix whose $(i,j)$th entry quantifies the similarity between covariates from nodes $i$ and $j$. NAC method incorporates node covariates into spectral clustering through the network-adjusted covariate matrix $\bm{A}\bm{X}+\bm{D}_{\alpha}\bm{X}\in\mathbb{R}^{N\times p}$ with diagonal weight matrix $\bm{D}_{\alpha}$ and covariate matrix $\bm{X}$. The SCORE method identifies communities through the ratios between the first leading eigenvector and the other leading eigenvectors.
For competitors, the true number of clusters from the simulated data and their default configurations are used during implementation. For GFL, we run our experiments with A100 GPU. Table 4 summarizes the number of model parameters and data points, when the node size is $N=120$, the latent dimension is $d=3$, and the sequence length is $n=100$. The neural network decoder consists of two hidden layers, each containing $32$ neurons.
Table 4: Number of parameters and data points.
Notation Description Quantity
$\bm{\phi}\$Neural network parameters $32\times 3+32\times 32+100\times 32\ +$ $32\times 1+32\times 1+100\times 1=4{,}484$
$\{\bm{\mu}_{i}\}_{i\in\mathcal{V}}$ Prior means $120\times 3=360$
$\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ Nodal series $120\times 100=12{,}000$
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ Graphical structures $|\mathcal{E}|$
B.1 Visualization
For both simulation and real data experiments, the choice of latent dimension $d=3$ permits visualizations of learned prior parameters in latent space. Figure 7 illustrates the estimated $\hat{\bm{\mu}}\in\mathbb{R}^{N\times 3}$ for a realization from Scenarios 1 - 3 with $d=3$. We can see that the estimated prior parameters have captured meaningful boundary between clusters, enforced by graph-fused LASSO regularization.
Figure 7: Illustration of estimated $\hat{\bm{\mu}}\in\mathbb{R}^{N\times d}$ for a realization from Scenarios 1 - 3, where the latent dimension is $d=3$. Each row, from top to bottom, corresponds to a scenario; each column, from left to right, displays a pairwise projection of the three latent dimensions. The node colors depict the ground truth labels.
Figure 8 displays the estimated $\hat{\bm{\mu}}\in\mathbb{R}^{58\times 3}$ from the California county temperature data analyzed in Section 4. From left to right, each panel displays a pairwise projection of the three latent dimensions. The first panel shows that counties within the same cluster are positioned closely in the latent space. The second and third panels of the graphic also show dome-shape relations for the third dimension (dim3) intersecting the first two dimensions (dim1 and dim2). Taken together, the three-dimensional latent space exhibits clear boundaries between clusters.
Figure 9 displays the estimated $\hat{\bm{\mu}}\in\mathbb{R}^{112\times 3}$ from the word co-occurrence network analyzed in Section 4. While the boundary between the blue and orange clusters are not visually apparent, the $k$-means algorithm and the silhouette score applied to the estimated prior parameters indicate the presence of two distinct clusters rather than a single one. In other words, partitioning the data points into two clusters results in more cohesive groupings with clearer separation. Overall, for the two real data experiments, the estimated prior parameters have captured meaningful separations between clusters in the latent space, driven by the graph-fused LASSO regularization.
Figure 8: Visualization of estimated $\hat{\bm{\mu}}\in\mathbb{R}^{58\times 3}$. Each data point represents one of the $58$ counties in California. Node colors represent the clusters detected by our method.
Figure 9: Visualization of estimated $\hat{\bm{\mu}}\in\mathbb{R}^{112\times 3}$. Each data point represents one of the $112$ words. Node colors represent the clusters detected by our method.
B.2 Sensitivity Analysis
The three simulation studies in Scenarios 1 - 3 are repeated for the proposed GFL method, with varying latent dimensions $d\in\{3,5,7,10\}$. Table 5 reports the results, with the values for $d=3$ directly duplicated from Tables 1, 2, and 3 for comparison. The proposed method is robust when the latent dimension is relatively low. However, when $d=10$, the performance declines significantly, likely due to over-fitting. In this case, the latent representation, which contains both temporal and structural information, could include excessive noise that deteriorates the clustering results.
Table 5: Sample means (one standard deviation) of our evaluation metrics for varied latent dimension $d\in\{3,5,7,10\}$ and Scenarios (Scen.) 1-3. The best metric is in bold.
Scen., NN $d$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ HOM $\uparrow$ COM $\uparrow$ PUR $\uparrow$
[…]
55 $\bm{98.80}\%(0.02)$ $\bm{99.17}\%(0.02)$ $\bm{99.72}\%(0.01)$ $\bm{98.79}\%(0.02)$ $\bm{98.80}\%(0.02)$ $\bm{99.72}\%(0.01)$
[…]
55 $\bm{99.40}\%(0.01)$ $\bm{99.62}\%(0.01)$ $\bm{99.88}\%(0.01)$ $\bm{99.40}\%(0.01)$ $\bm{99.40}\%(0.01)$ $\bm{99.88}\%(0.01)$
[…]
S2, $196$ 33 $\bm{99.71}\%(0.01)$ $\bm{99.75}\%(0.01)$ $\bm{99.92}\%(0.01)$ $\bm{99.71}\%(0.01)$ $\bm{99.71}\%(0.01)$ $\bm{99.92}\%(0.01)$
[…]
$10$ $88.59\%(0.11)$ $83.25\%(0.17)$ $87.48\%(0.12)$ $81.45\%(0.18)$ $\bm{99.93}\%(0.01)$ $\bm{99.98}\%(0.01)$
[…]
55 $\bm{99.47}\%(0.03)$ $\bm{99.22}\%(0.05)$ $\bm{99.42}\%(0.04)$ $\bm{99.17}\%(0.05)$ $\bm{99.96}\%(0.01)$ $\bm{99.99}\%(0.01)$
[…]